# Towards Context-Aware Domain Generalization: Understanding the Benefits and Limits of Marginal Transfer Learning

### v2

@@ Abstract
In this work, we analyze the conditions under which information about the context of an input $\mathbf{X}$ can improve the predictions of deep learning models in new domains. Following work in marginal transfer learning in Domain Generalization (DG), we formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same domain as the input itself. We offer a theoretical analysis of the conditions under which this approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice. Additionally, we contribute insights into the kind of distribution shifts for which the marginal transfer learning approach promises robustness. Empirical analysis shows that our criteria are effective in discerning both favorable and unfavorable scenarios. Finally, we demonstrate that we can reliably detect scenarios where a model is tasked with unwarranted extrapolation in out-of-distribution (OOD) domains, identifying potential failure cases. Consequently, we showcase a method to select between the most predictive and the most robust model, circumventing the well-known trade-off between predictive performance and robustness.
@@ 1 Introduction
Figure 1: Conceptual sketch of our setup and approach. A) Data-generating process (DGP) that fulfills our criteria. We assume that the environment $E$ is a source node that is not caused by any system variable and that the relationship between $\mathbf{X}$ and $\mathbf{Y}$ varies with the environment. $\mathbf{S}^{(n)}$ is a set of $n$ i.i.d. inputs available in the new environment (i.e. context). B) The workings of the context-aware model (aka the marginal transfer learning approach) in a test environment. A set-encoder generates a permutation-invariant representation $h(\mathbf{S}^{(n)})$ of the context. An inference network (e.g., a classifier) processes the representation along with the target input $\mathbf{X}$ and predicts the unknown outcome (e.g., label) of the target input. The set-representation can be combined with the input to reliably detect out-of-distribution (OOD) queries and prevent failure cases in domain generalization due to model misspecification.
Distribution shifts are the cause of many failure cases in machine learning [Hendrycks and Dietterich, 2019, Koh et al., 2021] and the root of various peculiar phenomena in classical statistics, such as Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021]. The Domain Generalization (DG) task seeks a model that is robust to distribution shifts by utilizing data from distinct environments during training [Muandet et al., 2013, Zhou et al., 2022].
@@
In the context of DG, marginal transfer learning enhances a model with context information to achieve better predictions Blanchard et al. [2021]. The “context” of a test instance is simply a set of samples that stems from the same environment as the instance itself and can be embedded, for instance, by permutation-invariant neural networks Edwards and Storkey [2017], Bloem-Reddy and Teh [2020]. In this work, we enhance the fundamental understanding of settings where marginal transfer learning in DG can reap benefits compared to baseline models.
@@
Consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$. Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands. It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$. This raises two questions: (i) Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$ and (ii) how should $E$ be represented to maximize the performance gain?
The first question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$ alone. For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region. Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically. A straightforward answer to the second question is to distinguish environments by discrete labels, but we propose to learn continuous embeddings instead. This allows us to compute representations for novel environments that enable teal-time adaptation and interpolation between the training environments. Moreover, discrete and continuous $E$ are equally informative for the known environments, ensuring no loss in information.
@@
Building on previous work in marginal transfer learning [Blanchard et al., 2021], we learn a continuous embedding of $E$ from auxiliary data using set encoders, as depicted in Figure 1. To systematically address the first question, we establish three criteria that delineate the circumstances in which $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and subsequently prove their necessity. Notably, two of these criteria are empirically testable using standard models and are shown to be necessary conditions for the success of the approach.
@@
When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and an increased risk of silent failures. While marginal transfer learning is not exempt from this “curse of extrapolation”, we find that it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [Müller et al., 2023]. Accordingly, we propose a method to select between models that are specialized in the in-distribution (ID) setting vs. models that are robust to out-of-distribution (OOD) scenarios on the fly. Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [Yang et al., 2022b, Müller et al., 2022, Magliacane et al., 2018].
In summary, our contributions are:
@@
  • We formalize the necessary and empirically verifiable conditions under which the marginal transfer learning approach can reap benefits from context information and improve on standard approaches;
  • We perform an extensive empirical evaluation and show that we can reliably detect failure cases when the necessary criteria of our theory are not met, or when extrapolation is required;
@@ 2 Method
In the following, we will briefly introduce context-aware models (aka marginal transfer learning) in DG. We then discuss criteria under which we can expect to reap benefits from the additional set-representation. Afterwards, we explain the theoretical data-generating process that matches these criteria, providing insight into the distribution shifts for which context-aware approaches may prove advantageous in practice. Finally, we discuss the process of identifying new environments that demand extrapolation, potentially leading to failure cases.
@@ 2.1 Notation
We denote inputs $\mathbf{X}\in\mathcal{X}$ and outputs as $\mathbf{Y}\in\mathcal{Y}$, without any strict requirements on the input and output spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ a set of $n$ i.i.d. samples (i.e., a set of further inputs) from the given domain. The domain label $E$ is only known during training time and unknown during inference.
@@ 2.2 Context-Aware Models
A context-aware model consists of two key components (also illustrated in Figure 1): (i) a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set encoder”) with parameters $\psi$ that maps a set input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and (ii) an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction. The complete model is denoted as $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})=f_{\boldsymbol{\phi}}(\mathbf{X},h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)}))$ with parameters $\boldsymbol{\theta}=(\boldsymbol{\psi},\boldsymbol{\phi})$ for short. For a given supervised learning task, we aim to find the minimum to the following optimization problem
@@ 2.3 Criteria for Improvement
In the following, we establish criteria under which context information allows to exploit the distribution shifts between environments and yield improved predictions. In total, we propose three criteria that are necessary to achieve incremental improvement. In Theorem 2.1, we show how these criteria are related to each other. In the formulations below, $I(\mathbf{X};\mathbf{Y})$ denotes the mutual information between random vectors $\mathbf{X}$ and $\mathbf{Y}$ and $I(\mathbf{X};\mathbf{Y}\mid\mathbf{Z})$ denotes the conditional mutual information given a third random vector $\mathbf{Z}$. The symbol $\perp$ (resp. ⟂$\not\perp$̸) between two random vectors $\mathbf{X}$ and $\mathbf{Y}$ is used to express that the random vectors are independent (resp. dependent) or conditionally independent (resp. dependent) given a third random vector $\mathbf{Z}$.
@@
In Figure 1, an instance $\mathbf{X}$ cannot be assigned with complete certainty to an environment. Consequentially, further data provides additional information about the environment. In general, the more data we consider, the better we can predict the originating environment. Crucially, this criterion is not satisfied, if we can recover the origin environment from the singleton input $\mathbf{X}$ alone.
@@
This criterion can serve as a sanity check in case we have an oracle that can identify the origin environment of the data with perfect accuracy.
In what follows, we show that Criterion 2.2 and Criterion 2.3 are necessary conditions for Criterion 2.1. We furthermore prove that if we can extract the environment label fully from $\mathbf{S}^{(n)}$, then Criterion 2.2 and Criterion 2.3 are sufficient conditions for Criterion 2.1. We generalize this result for the case where the environment label is not inferable with 100% accuracy in Appendix C.
@@
The proof for this theorem can be found in subsection C.3. Unfortunately, we cannot conclude that $Y\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general. A counterexample where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in subsection C.2. It is also worth noting that Criterion 2.3 might be achievable while Criterion 2.2 is unattainable and vice versa. For instance, when we can infer the originating environment from one sample (Criterion 2.2 is not attainable), the relation between $\mathbf{X}$ and $\mathbf{Y}$ might still vary with the environment (Criterion 2.3 is achievable).
@@ 2.4 Source Component Shift
Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satifsied. Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [Quinonero-Candela et al., 2008]. The source component shift can be described by the graphical model in Figure 1, where the environment directly affects both the input $\mathbf{X}$ and the outcome $\mathbf{Y}$. Problems that conform to the graph in Figure 1 have two important implications. First, the input distribution changes whenever the environment changes. [^1] In our case, we need a stronger assumption: the domains cannot be deduced from a single input, as defined in Criterion 2.2. Second, the relationship between inputs and outcomes varies with the environment (corresponding to Criterion 2.1). For more details on this kind of distribution shift, we refer the reader to Quinonero-Candela et al. [2008]. It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021], which supplies a proof-of-concept for our approach (see Experiment 1). An important point to highlight is that the frequently encountered covariate shift where only $P(\mathbf{X})$ in $P(\mathbf{X},\mathbf{Y})=P(\mathbf{Y}\mid\mathbf{X})P(\mathbf{X})$ varies between environments [Quinonero-Candela et al., 2008], does not conform to the conditions specified in Criterion 2.3. Hence, context-aware models do not provide advantages when compared to standard models under covariate shift (under the assumption all models converged to the optimum).
@@ 2.5 Detection of Novel Environments
During test time, data could either originate from an environment that corresponds to one of the training environments (but its origins are unknown) or from a previously unseen environment. In the following, we explain how we aim to detect the second case that might result in potential failure cases due to fundamental challenges in extrapolation. Following Müller et al. [2023], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$. As a score function, we consider the distance of $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to the $k$-nearest neighbors in the training data in the feature space of the set-encoder. Accordingly, set-representations that elicit a score surpassing a certain threshold are considered to originate from a novel environment.
Following the approach in Müller et al. [2023], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment. To establish this threshold, we consider the $q$-th percentile of scores obtained from the validation set. We also compare our novel environment detector with the same score function computed from singleton features $g(\mathbf{X})$ alone (see Table 1 for a preview).
@@ 3.1 Domain Generalization
The aim of Domain Generalization (DG) is to train models that generalize well under distribution shifts [Muandet et al., 2013, Zhou et al., 2022]. In contrast to domain adaptation (DA) [Wang and Deng, 2018], where samples from the test domain are given during training, in DG no knowledge about the test environment is available during training. Both, DG and DA, involve access to data from multiple domains during training exploiting the heterogeneity between domains. For an overview on DG see for instance Wang et al. [2022], Zhou et al. [2022].
As a middle-ground between DA and DG, test-time adaptation (TTA) [Liang et al., 2023] and marginal transfer learning [Blanchard et al., 2021] involves the provision of unlabeled samples during test time. In TTA, the model is typically fine-tuned according to the provided samples [Liang et al., 2023]. Marginal transfer learning in DG refers to the setting where access to the marginal feature distribution is given during testing and provided to the model via $\frac{1}{n}\sum_{i=1}^{n}\sigma(\mathbf{X}_{i})$ [Blanchard et al., 2021]. In this context, $\sigma$ represents a feature extraction model applied to each $\mathbf{X}_{i}$, and the summation ensures that the operation remains invariant under permutations. Different choices for $\sigma$ are explored in the literature. Specifically, Zhang et al. [2021] utilizes CNNs, Blanchard et al. [2021], Dubey et al. [2021] incorporates kernel embeddings, and Bao and Karaletsos [2023] adopts patch embeddings. While these works highlight the advantages of their methodologies, they fall short in analyzing the exact settings and conditions that contribute to these benefits compared to a standard model. Additionally, they do not address the detection of potential failure cases or the utilization of context embedding for model selection. A recent method proposes to use transformers to exploit previously seen samples instead of learning permutation-invariant embedding [Gupta et al., 2023]
The trade-off between ID performance and OOD performance is well-known [Müller et al., 2022, Magliacane et al., 2018, Zhang et al., 2023]. A method to mitigate this trade-off was proposed in Zhang et al. [2023]. Their approach presupposes knowledge of the data’s originating environment, whereas our goal is to infer it.

### v3

# Reliable ML From Unreliable Data

@@ Abstract
In this work, we analyze the conditions under which information about the context of an input data point can improve the predictions of deep learning models in new domains. Following work in marginal transfer learning and domain generalization, we formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same domain as the input itself. We offer a theoretical analysis of the conditions under which this approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice. Additionally, we contribute insights into the kind of distribution shifts for which the marginal transfer learning approach promises robustness. Empirical analysis shows that our criteria are effective in discerning both favorable and unfavorable scenarios. Finally, we demonstrate that we can reliably detect scenarios where a model is tasked with unwarranted extrapolation in out-of-distribution (OOD) domains, identifying potential failure cases. Consequently, we showcase a method to select between the most predictive and the most robust model, circumventing the well-known trade-off between predictive performance and robustness.
@@ 1 Introduction
Distribution shifts are the cause of many failure cases in machine learning [1, 2] and the root of various peculiar phenomena in classical statistics, such as Simpson’s paradox [3, 4]. Domain Generalization (DG) seeks models that are robust to distribution shifts by utilizing data from distinct environments during training [5, 6]. In the context of DG, marginal transfer learning enhances a model with context information to achieve better predictions [7]. The “context” of a test instance is a set of samples that stems from the same environment as the instance itself and can be embedded, for instance, by permutation-invariant neural networks [8]. In this work, we enhance the fundamental understanding of settings where marginal transfer learning in DG can reap benefits compared to baseline models.
@@
Consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$. Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands. It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$. This raises a key question: Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$? This question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$ alone. For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region. Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically.
Figure 1: Conceptual sketch of our setup and approach. A) Data-generating process (DGP) that fulfills our criteria. We assume that the environment $E$ is a source node that is not caused by any system variable and that the relationship between $\mathbf{X}$ and $\mathbf{Y}$ varies with the environment. $\mathbf{S}^{(n)}$ is a set of $n$ i.i.d. inputs available in the new environment. The bidirectional arrow indicates that the causal relation between $\mathbf{X}$ and $\mathbf{Y}$ could be explained by a common cause or $\mathbf{Y}$ causing $\mathbf{X}$ (or vice versa). B) The context-aware model (marginal transfer learning approach) in a test environment. A set-encoder generates a permutation-invariant representation $h(\mathbf{S}^{(n)})$ of the context. An inference network processes the representation along with the target input $\mathbf{X}$ and predicts the unknown outcome of the target input. The set-representation can be combined with the input to reliably detect out-of-distribution queries and prevent failure cases in domain generalization due to model misspecification.
@@
Building on previous work in marginal transfer learning [7], we aim to learn a continuous embedding of $E$ from auxiliary data using set-encoders, as depicted in Figure 1. We then establish three criteria that delineate the circumstances in which $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and subsequently prove their necessity. Notably, two of these criteria are empirically testable using standard models and are shown to be necessary conditions for the success of the approach.
@@
When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and an increased risk of silent failures. While marginal transfer learning is not exempt from this “curse of extrapolation”, we find that it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [9]. Accordingly, we propose a method to select between models that are specialized in the ID setting versus models that are robust to OOD scenarios on the fly. Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [10, 11, 12]. In summary, our contributions are:
@@
  • We formalize the necessary and empirically verifiable conditions under which the marginal transfer learning can improve on standard approaches;
  • We show empirically that we can identify cases where context-aware models offer no advantages or when dangerous extrapolation is necessary;
@@ 2.1 Notation
We denote inputs $\mathbf{X}\in\mathcal{X}$ and outputs as $\mathbf{Y}\in\mathcal{Y}$, without any strict requirements on the input and output spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ a set of $n$ further i.i.d. samples from a given domain, whose label $E$ is only known during training time.
@@ 2.2 Context-Aware Models
A context-aware model consists of two key components (also illustrated in Figure 1): (i) a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set-encoder”) with parameters $\psi$ that maps a set-input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and (ii) an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction. The complete model is denoted as $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})=f_{\boldsymbol{\phi}}(\mathbf{X},h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)}))$ with parameters $\boldsymbol{\theta}=(\boldsymbol{\psi},\boldsymbol{\phi})$ for short. For a given supervised learning task, we consider the optimization problem
@@ 2.3 Criteria for Improvement
In the following, we establish criteria under which context information allows to exploit the distribution shifts between environments and yield improved predictions.
In total, we propose three criteria that are necessary to achieve incremental improvement. In Theorem 2.1, we show how these criteria are related to each other. In the formulations below, $I(\mathbf{X};\mathbf{Y})$ denotes the mutual information between random vectors $\mathbf{X}$ and $\mathbf{Y}$ and $I(\mathbf{X};\mathbf{Y}\mid\mathbf{Z})$ denotes the conditional mutual information given a third random vector $\mathbf{Z}$. The symbol $\perp$ (resp. $\not\perp$) between two random vectors $\mathbf{X}$ and $\mathbf{Y}$ is used to express that the random vectors are independent (resp. dependent) or conditionally independent (resp. dependent) given a third random vector $\mathbf{Z}$.
@@
In Figure 2, an instance $\mathbf{X}$ cannot be assigned with complete certainty to an environment. Consequentially, further data provides additional information about the environment. In general, the more data we consider, the better we can predict the originating environment. Crucially, this criterion is not satisfied if we can recover the origin environment from the singleton input $\mathbf{X}$ alone.
@@
This criterion can serve as a sanity check in case we have an oracle that can identify the origin environment of the data with perfect accuracy. In what follows, we show that Criterion 2.2 and Criterion 2.3 are necessary conditions for Criterion 2.1. We furthermore prove that if we can extract the environment label fully from $\mathbf{S}^{(n)}$, then Criterion 2.2 and Criterion 2.3 are sufficient conditions for Criterion 2.1.
@@
In our experiments, we observe that a function $g(\mathbf{S}^{(n)})=E$ can already be found for small $n$ (see for instance Figure 2). In this case, we obtain $I(\mathbf{S}^{(n)};\,\mathbf{Y}\mid\mathbf{X})=I(E;\,\mathbf{Y}\mid\mathbf{X})$ and Criterion 2.2 and Criterion 2.3 are sufficient to obtain Criterion 2.1. Unfortunately, we cannot conclude that $Y\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general. A counterexample where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in Section C.2. We furthermore provide the proof of the theorem, an illustration for the theorem as well as a generalization of (c) in Appendix C.
It is worth noting that model misspecification adds another layer of uncertainty when verifying the criteria. In cases where determining the correct mutual information is not feasible (for instance, when $p(\mathbf{Y}\mid\mathbf{X})$, $p(\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}$), or $p(\mathbf{Y}\mid\mathbf{X},E)$ cannot be learned adequately), two primary issues may emerge. Firstly, the effective utilization of the set-input $\mathbf{S}^{(n)}$ (or E) may be hindered due to either the restricted expressive power of the model class or a scarcity of training data. As a result, the context-aware model might not improve on a baseline model that only utilizes $\mathbf{X}$. Consequentially the criteria could seem unattainable while they actually are. Secondly, after training, we might observe an apparent advantage of the approximation of $p(\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)})$ or $p(\mathbf{Y}\mid\mathbf{X},E)$ over the approximation of $p(\mathbf{Y}\mid\mathbf{X})$, despite the true model not conferring any advantage. In this scenario, the criteria may appear to be satisfied, whereas in reality they are not. An example of this case can be easily constructed by considering a non-linear $p(\mathbf{Y}\mid\mathbf{X})$ and a linear function class.
@@ 2.4 Source Component Shift
Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satisfied. Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [13]. The source component shift can be described by the graphical model in Figure 1, where the environment directly affects both the input $\mathbf{X}$ and the outcome $\mathbf{Y}$. Problems that conform to the graph in Figure 1 have two important implications. First, the input distribution changes whenever the environment changes. Second, the relationship between inputs and outcomes varies with the environment (corresponding to Criterion 2.1). For more details on this kind of distribution shift, we refer the reader to [13]. It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [3, 4], which supplies a proof-of-concept for our approach (see Experiment 1). An important point to highlight is that the frequently encountered covariate shift where only $P(\mathbf{X})$ in $P(\mathbf{X},\mathbf{Y})=P(\mathbf{Y}\mid\mathbf{X})P(\mathbf{X})$ varies between environments [13], does not conform to the conditions specified in Criterion 2.3. Hence, context-aware models do not provide advantages when compared to standard models under covariate shift.
@@ 2.5 Detection of Novel Environments
During test time, data could either originate from an environment that corresponds to one of the training environments (but its origins are unknown) or from a previously unseen environment. In the following, we explain how we aim to detect the second case that might result in potential failure cases due to fundamental challenges in extrapolation. Following [9], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$. As a score function, we consider the distance of $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to the $k$-nearest neighbors in the training data in the feature space of the set-encoder. Accordingly, set-representations that elicit a score surpassing a certain threshold are considered to originate from a novel environment.
Following the approach in [9], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment. To establish this threshold, we consider the $q$-th percentile of scores obtained from the validation set. We also compare our novel environment detector with the same score function computed from singleton features $g(\mathbf{X})$ alone (see Table 2 for a preview).
@@ 3.1 Domain Generalization
Domain Generalization (DG) trains models to perform under distribution shifts without access to test environments [5, 6]. In contrast, Domain Adaptation (DA) assumes test samples are available during training [14]. Both exploit multiple source domains, but DG is strictly test-agnostic. Non-marginal DG approaches fall into three groups [15]: data manipulation [16, 17], robust representation learning [18, 19], and learning strategy modification [20, 21] (see 15, 6 for reviews).
Between DA and DG lie test-time adaptation (TTA) and marginal transfer learning. TTA adapts to unlabeled test samples, often via fine-tuning or domain metadata [22, 23]. Marginal transfer instead assumes access to the marginal feature distribution $\frac{1}{n}\sum_{i}\sigma(\mathbf{X}_{i})$ [7], with $\sigma$ implemented via CNNs [24], kernel embeddings [7, 25], or patch embeddings [26]. While Blanchard et al. [7] analyze kernel embeddings theoretically, existing work leaves open conditions for effectiveness, failure detection, and context-aware model selection. A recent alternative replaces permutation-invariant embeddings with transformers that exploit sample order [27].
Marginal transfer parallels in-context learning: labeled samples define context in the latter [28, 29], while unlabeled samples do so in the former. Finally, balancing in-domain and out-of-domain performance remains a central challenge [11, 12, 30]. Methods like Zhang et al. [30] mitigate this trade-off when domain identity is known, whereas our goal is to infer it.